Sol Summary – February 9th

Sol: Sol #5
Summary Title: Snow at MDRS
Author’s name: Annalea Beattie
Mission Status: Nominal
Sol Activity Summary: All field work is postponed today due to a layer of snow covering the desert. We regrouped after breakfast and met to set the goals for the day. Because of the weather forecast, Bharti and Clare spent some time planning the minimum schedule of EVAs necessary to complete everyone’s goals for field work in the following week. Bharti also read and took notes on the geomorphology of this area. She then prepared a base map. Aditya helped Mehnaz with evapotranspiration calculations for her research and then he had his scheduled shower. Later he continued his analysis of star images, specifically in response to the feedback given by Director of Observatories, Peter Detterline, and in relation to his image of NGC5904. Rajvi performed her engineer jobs and then discussed and examined variations of concretions with Clare. To minimize tools and save time, she is learning how to use her watch to mark points and trace her steps in the field. She also cooked a delicious lunch. Daniel spent much of the day in the Science Dome. In the morning he made two 5ml dilutions of the 5 different types of liquid media he brought with him. These were inoculated with 0.25g of the two different soil samples collected yesterday on EVA 4 and placed in the incubator which is set at 37C/99F to see if anything will grow. After lunch Daniel baked bread with Bharti and then extracted DNA from yesterday’s samples from EVA 4 in the Science Dome. I reorganized my notes for our upcoming outreach event with Women in Space, South Asia, did yoga and baked more bread. After two pm, I used the table and the magnification lamp in the Science dome to do several hours drawing and painting of concretion samples. Mehnaz gardened in the Green Hab, harvested greens for dinner and worked on the equations for her research.
Look Ahead Plan: Today while the snow melted we discussed our postponed EVAS and what might be possible for tomorrow, considering the forecast for snow, rain, thunderstorms and other hazardous weather predicted in the next few days. The forecast is not good for our fieldwork but we are trying to think about what is do-able and whether we can work close to the hab, weather permitting. Meanwhile we will ensure all our research is ready to go, as soon as the weather clears.
Anomalies in work: None
Weather: Snow covering the desert and sunny during the day with a high of 41F. Rain and cold with a wind chill factor of 30F and 22F overnight tonight with a chance of new snow accumulation of half an inch.
Crew Physical Status: Nominal. The crew are working well together as a team, everyone is healthy and morale is good.
EVA: None
Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, Astronomy Report, EVA request.
Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site